# OSCAR: One-Step Diffusion Codec for Image Compression Across Multiple Bit-rates

### v3

@@ Abstract
Pretrained latent diffusion models have shown strong potential for lossy image compression, owing to their powerful generative priors. Most existing diffusion-based methods reconstruct images by iteratively denoising from random noise, guided by compressed latent representations. While these approaches have achieved high reconstruction quality, their multi-step sampling process incurs substantial computational overhead. Moreover, they typically require training separate models for different compression bit-rates, leading to significant training and storage costs. To address these challenges, we propose a one-step diffusion codec across multiple bit-rates. termed OSCAR. Specifically, our method views compressed latents as noisy variants of the original latents, where the level of distortion depends on the bit-rate. This perspective allows them to be modeled as intermediate states along a diffusion trajectory. By establishing a mapping from the compression bit-rate to a pseudo diffusion timestep, we condition a single generative model to support reconstructions at multiple bit-rates. Meanwhile, we argue that the compressed latents retain rich structural information, thereby making one-step denoising feasible. Thus, OSCAR replaces iterative sampling with a single denoising pass, significantly improving inference efficiency. Extensive experiments demonstrate that OSCAR achieves superior performance in both quantitative and visual quality metrics. The code and models will be released at https://github.com/jp-guo/OSCAR.
@@ 1 Introduction
With the explosive growth of multimedia content, the need for efficient transmission and storage of high-resolution images has become increasingly critical. Traditional image compression has long relied on manually crafted codecs [50, 6, 10], which tend to falter at very low bit rates ($\leq 0.2$ bpp), where blocking and ringing artifacts severely degrade perceptual quality. Recent research has shifted towards learning-based transform-coding systems [5, 13, 21] that jointly optimize an end-to-end rate–distortion (R–D) objective [42]. Although these neural codecs generally outperform the hand-crafted compression algorithms, the focus on minimizing R–D inevitably sacrifices fine-grained realism at extreme compression bit-rates [8], producing images that appear synthetic or over-smoothed. In addition, supporting multiple compression bit-rates generally means training multiple models, each optimized with a rate–distortion loss weighted by a different Lagrange multiplier ($\lambda$). Maintaining one model per $\lambda$ dramatically inflates both training time and storage requirements.
@@
Recent breakthroughs in diffusion models [23, 43], most notably large-scale text-to-image latent diffusion models (LDMs) [39], have revealed powerful generative priors that can be repurposed for a wide range of downstream vision tasks [56, 9, 19]. Building on this insight, diffusion-based codecs [44, 30, 32] have been proposed for image compression. These models typically perform quantization in the latent space to enable compression, achieving lower distortion at a given rate than conventional CNN- or Transformer-based methods. Some works [11] further integrate pretrained LDMs with a vector-quantization mechanism [48, 16], allowing the compression rate to be predetermined via a fixed hyper-encoder. Their practicality, however, is still hindered by two issues: (i) the multiple denoising iterations are required at inference time, which impose substantial inference computational overhead; and (ii) achieving multiple bit-rates still requires training separate diffusion networks. Since diffusion models typically have a larger model size than conventional learned codecs, this one-model-per-rate strategy further amplifies both training and storage costs. These constraints limit real-world deployment, particularly in resource-constrained scenarios.
Figure 1: Qualitative comparison between HiFiC [35], MS-ILLM [37], CDC [54], PerCo [11], and our OSCAR. OSCAR is capable of reconstructing complex textures with high realism. OSCAR effectively reconstructs complex textures with high realism. In contrast, HiFiC, MS-ILLM, and PerCo exhibit noticeable loss of fine structures and introduce visible artifacts, while CDC suffers from significant color shifts. Please zoom in for a better view.
@@
To address the aforementioned challenges, we propose OSCAR, a one-step diffusion codec across multiple bit-rates. Unlike previous approaches [54, 32, 11] that start from Gaussian noise, we introduce the concept of pseudo diffusion timestep, bridging the quantization process and the forward diffusion. This idea is inspired by a classical signal processing result that quantization noise, especially at higher bit-rates, can be approximated as additive Gaussian noise [7]. Formally, we posit that the quantized latents $\tilde{\mathbf{z}}$ can be decomposed as $\tilde{\mathbf{z}}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}$, where $\boldsymbol{\epsilon}$ is approximately Gaussian noise and statistically independent of $\mathbf{z}_{0}$, $\bar{\alpha}_{t}$ is the diffusion noise schedule, and $t$ is the pseudo diffusion timestep that depends on the bit-rate. Under the assumption that $\boldsymbol{\epsilon}$ is orthogonal to $\mathbf{z}_{0}$ in expectation, the expected cosine similarity between $\tilde{\mathbf{z}}$ and $\mathbf{z}_{0}$ satisfies $\mathbb{E}[\operatorname{sim(\tilde{\mathbf{z}},\mathbf{z}_{0})}]\approx\sqrt{\bar{\alpha}_{t}}$ (Please refer to Appendix for a full derivation). To make this relationship more consistent, we adopt a representation alignment training between $\tilde{\mathbf{z}}$ and $\mathbf{z}_{0}$. We empirically observe that this training paradigm enables their cosine similarity to converge to a fixed value for each bit-rate. Thus, by matching measured cosine similarities to the theoretical form, each bit-rate $r$ can be directly mapped to a pseudo diffusion timestep $t$. Our experiments provide strong support for this modeling, as the measured distribution of $\boldsymbol{\epsilon}$ closely follows the assumed Gaussian distribution.
@@
Figure 2: Left: Our method bridges quantization and diffusion by mapping bit-rates to pseudo diffusion timesteps. Right: BD-rates measured by DISTS [15], with circle radius indicating multiply-accumulate operations (MACs). OSCAR achieves a favorable quality-efficiency tradeoff.
@@ 2.1 Neural Image Compression
Deep learning has enabled neural image compression methods to outperform traditional codecs such as JPEG [50], BPG [6], and VVC [10]. Pioneering work [4] introduced end-to-end autoencoders for learned compression, followed by efforts  [5, 36, 13, 21] to jointly optimize rate and distortion through improved latent representations. Yet, distortion-based losses (e.g., MSE) often lead to distribution shifts between reconstructed and natural images [8], motivating research [3, 22, 1] to enhance perceptual quality. HiFiC [35] explores the integration of GANs [18] with compression. MS-ILLM [37] improves reconstruction quality through improved discriminator design. Other methods leverage textual semantics [28], causal modeling [20], or decoding-time guidance [1] to further boost perceptual fidelity. However, the aforementioned methods typically require training separate models for each bit-rate, which results in substantial overhead.
To address the inefficiency of training separate models for each bit-rate, recent research has explored adaptive compression strategies that enable flexible bit-rate control. In particular, progressive compression  [47, 34, 55] has attracted significant interest for its ability to produce scalable bitstreams. DPICT [29] encodes latent features into trit-plane bitstreams, allowing fine-grained scalability. CTC [25] extends this idea by introducing context-based trit-plane coding. Meanwhile, variable-rate approaches  [12, 51] have demonstrated effectiveness by dynamically adjusting scalar parameters or employing conditional convolutions to accommodate diverse quality levels. Despite their flexibility, these methods tend to suffer from notable quality degradation at very low bit-rates.
@@ 2.2 Diffusion Models
Diffusion models are powerful generative frameworks that transform random noise into coherent data through a multi-step denoising process [23]. Denoising Diffusion Implicit Models (DDIM) [43] accelerate this process without compromising quality, influencing many subsequent advances. Latent Diffusion Models (LDMs) further improve efficiency by operating in a compressed latent space. Stable Diffusion (SD) [39] advances the LDM architecture into a scalable text-to-image system, enabling flexible multimodal conditioning while maintaining strong image fidelity. In our work, we build on SD to benefit from its strong image generation capabilities.
Beyond image synthesis, diffusion models have also been explored in image compression  [41, 30]. Early work [23] explored diffusion-based compression through reverse channel coding, with an emphasis on rate-distortion trade-offs. Follow-up studies [44] extended this line by applying the technique to lossy transmission of Gaussian samples. Recent approaches further leverage large-scale pretrained diffusion models to enhance performance [49]. CDC [54] replaces the VAE decoder with a conditional diffusion model, reconstructing images via reverse diffusion conditioned on a learned content latent. PerCo [11] adopts a VQ-VAE-style codebook for discrete representation and incorporates global textual descriptions to guide the iterative decoding process. DiffEIC [32] uses a lightweight control module to guide a frozen diffusion model with compressed content. While these models achieve strong results, their multi-step denoising process remains computationally expensive. Improving sampling efficiency is therefore essential for practical diffusion-based image compression.
@@ 3.1 Preliminaries: One-Step Latent Diffusion Model
Latent Diffusion Models (LDMs) [39] operate in a lower-dimensional latent space rather than pixel space, enabling more efficient training and generation. The forward diffusion process perturbs a clean latent variable $\mathbf{z}$ by injecting Gaussian noise. At timestep $t$, the corrupted latent is given by:
@@
where the noise schedule is encoded as $\bar{\alpha}_{t}=\prod_{i=1}^{t}(1-\beta_{i})$, and $\beta_{i}$ is the variance. The reverse process aims to progressively recover the original latent through a learned denoising network [23, 43], typically implemented as a U-Net [40] structure in LDMs. Standard diffusion models perform multi-step denoising via a Markov chain, where each reverse step is modeled as:
@@ 3.3 Stage 1: Learning Quantization Hyper-Encoders
Figure 4: Distribution and QQ-plot of residual noise $\boldsymbol{\epsilon}$ in Eq. (11) before and after representation alignment. Top: red line is the standard Gaussian distribution; blue histogram is the measured distribution. Bottom: red line is the theoretical quantile line; blue dots are the measured values.
Bit-rate-specific hyper-encoders. While latent features benefit from low spatial resolution and are well-suited for compression, they still require significant storage, as they encapsulate high-dimensional semantic information [39, 38]. To address this, we apply hyper-encoders to quantize the latent representations. Given an input image, OSCAR first uses the Stable Diffusion (SD) VAE encoder to extract the latent representation $\mathbf{z}_{0}$. This step reduces spatial resolution (e.g., a 1024$\times$1024 image yields a latent of shape 4$\times$128$\times$128). To further compress $\mathbf{z}_{0}\in\mathbb{R}^{4\times h\times w}$, we employ a lightweight hyper-encoder that downsamples and quantizes it into $\mathbf{z}_{q}\in\mathbb{R}^{M\times h^{\prime}\times w^{\prime}}$, where $M$ is the dimension of the quantization space. Specifically for the architecture, the front-end of the hyper-encoder consists of several residual blocks, followed by a vector quantization (VQ) codebook. With downsampling ratio $s$ and codebook size $V$, the resulting bit-rate is $bpp=\log_{2}V/64s^{2}$. For decoding, the back-end module of the hyper-encoder upsamples $\mathbf{z}_{q}$ and projects it back to the original latent space, yielding the quantized latent feature $\tilde{\mathbf{z}}\in\mathbb{R}^{4\times h\times w}$. The back-end module employs attention layers and additional residual blocks to enhance reconstruction quality. We provide the detailed architecture in the Appendix.
@@
where N is the number of vectors in $\mathbf{z}_{0}$ and $r$ is the desired compression bit-rate. As in VQ-VAE [48], since vector-quantization (VQ) operation is non-differentiable, the VQ loss is defined as:
@@
where $\operatorname{sg}$ is the stop-gradient operation, and $\mathbf{z}_{q}$ is the mapping of $\mathbf{z}_{0}$ to its nearest codebook entry. Following the method in [16], we adopt an exponential moving average (EMA) update for the codebook, which enhances training stability and eliminates the need for the first term in Eq. (11). Thus, the overall representation alignment objective is:
@@
We argue that representation alignment is critical for modeling compressed latents as intermediate diffusion states (We will elaborate on this in Section 3.4). To support this, we compare the measured additive noise before and after training. As shown in Fig. 4, prior to training, the residual noise exhibits no clear structure and resembles a uniform distribution within the range of 2.5 to 2.5. The corresponding quantile-quantile plots (QQ-plots) [53] deviate substantially from the theoretical Gaussian line. This indicates that the latents prior to alignment contain only unstructured noise. In contrast, after training, the residual noise closely follows a Gaussian distribution, and the QQ-plots align well with the theoretical line, indicating that the residuals become statistically well-behaved.
@@ 3.4 Stage 2: Unified Training for One-Step Reconstruction
Specifically, given the predefined bit-rate set $\mathcal{R}$, we fine-tune all the corresponding hyper-encoders and apply LoRA [24] to update the diffusion model. The overall loss combines a representation alignment loss $\mathcal{L}_{\text{repa}}^{r}$, a perceptual loss $\mathcal{L}_{\text{per}}$, and an adversarial loss $\mathcal{L}_{\mathcal{G}}$:
@@
The alignment loss $\mathcal{L}_{\text{repa}}$ preserves structural consistency between the quantized and original latents. The perceptual loss $\mathcal{L}_{\text{per}}$ includes MSE and DISTS [15] terms. Following the practice of [14], we additionally use the Sobel operator to extract the edge of the image and measure the corresponding edge-aware DISTS loss:
$\mathcal{L}_{\text{per}}(\phi,\theta)\coloneqq\mathbb{E}_{\mathbf{I}\sim\mathcal{D}}\left[\mathcal{L}_{2}(\mathbf{I},\hat{\mathbf{I}})+\mathcal{L}_{\text{DISTS}}(\mathbf{I},\hat{\mathbf{I}})+\mathcal{L}_{\text{EA-DISTS}}(\mathcal{S}(\mathbf{I}),\mathcal{S}(\hat{\mathbf{I}}))\right],$ (15)
where $\hat{\mathbf{I}}=\mathcal{G}_{\phi,\theta}(\mathbf{z}_{0})$, $\mathcal{S}(\cdot)$ is the Sobel operator. Finally, we introduce a discriminator $\mathcal{D}_{\psi}$ and adopt a GAN loss [3] in latent space to encourage realism of the reconstructed image:

### v6

# OSCAR: One-Step Diffusion Codec Across Multiple Bit-rates

@@ Abstract
Pretrained latent diffusion models have shown strong potential for lossy image compression, owing to their powerful generative priors. Most existing diffusion-based methods reconstruct images by iteratively denoising from random noise, guided by compressed latent representations. While these approaches have achieved high reconstruction quality, their multi-step sampling process incurs substantial computational overhead. Moreover, they typically require training separate models for different compression bit-rates, leading to significant training and storage costs. To address these challenges, we propose a one-step diffusion codec across multiple bit-rates. termed OSCAR. Specifically, our method views compressed latents as noisy variants of the original latents, where the level of distortion depends on the bit-rate. This perspective allows them to be modeled as intermediate states along a diffusion trajectory. By establishing a mapping from the compression bit-rate to a pseudo diffusion timestep, we condition a single generative model to support reconstructions at multiple bit-rates. Meanwhile, we argue that the compressed latents retain rich structural information, thereby making one-step denoising feasible. Thus, OSCAR replaces iterative sampling with a single denoising pass, significantly improving inference efficiency. Extensive experiments demonstrate that OSCAR achieves superior performance in both quantitative and visual quality metrics. The code and models are available at https://github.com/jp-guo/OSCAR.
@@ 1 Introduction
With the explosive growth of multimedia content, the need for efficient transmission and storage of high-resolution images has become increasingly critical. Traditional image compression has long relied on manually crafted codecs wallace1991jpeg ; bpg ; bross2021overview , which tend to falter at very low bit rates ($\leq 0.2$ bpp), where blocking and ringing artifacts severely degrade perceptual quality. Recent research has shifted towards learning-based transform-coding systems balle2018variational ; cheng2020learned ; he2022elic that jointly optimize an end-to-end rate–distortion (R–D) objective shannon1948mathematical . Although these neural codecs generally outperform the hand-crafted compression algorithms, the focus on minimizing R–D inevitably sacrifices fine-grained realism at extreme compression bit-rates blau2019rethinking , producing images that appear synthetic or over-smoothed. In addition, supporting multiple compression bit-rates generally means training multiple models, each optimized with a rate–distortion loss weighted by a different Lagrange multiplier ($\lambda$). Maintaining one model per $\lambda$ dramatically inflates both training time and storage requirements.
@@
Recent breakthroughs in diffusion models sohl2015deep ; ho2020denoising ; song2020denoising , most notably large-scale text-to-image latent diffusion models (LDMs) rombach2022high , have revealed powerful generative priors that can be repurposed for a wide range of downstream vision tasks bansal2022cold ; brooks2023instructpix2pix ; guo2025compression . Building on this insight, diffusion-based codecs Theis2022b ; lei2023textsketch ; li2024towards have been proposed for image compression. These models typically perform quantization in the latent space to enable compression, achieving lower distortion at a given rate than conventional CNN- or Transformer-based methods. Some works careil2024towards further integrate pretrained LDMs with a vector-quantization mechanism oord2017neural ; esser2021taming , allowing the compression rate to be predetermined via a fixed hyper-encoder. Their practicality, however, is still hindered by two issues: (i) the multiple denoising iterations are required at inference time, which impose substantial inference computational overhead; and (ii) achieving multiple bit-rates still requires training separate diffusion networks. Since diffusion models typically have a larger model size than conventional learned codecs, this one-model-per-rate strategy further amplifies both training and storage costs. These constraints limit real-world deployment, particularly in resource-constrained scenarios.
Figure 1: Qualitative comparison. OSCAR is capable of reconstructing complex textures with high realism. OSCAR effectively reconstructs complex textures with high realism.
@@
To address the aforementioned challenges, we propose OSCAR, a one-step diffusion codec across multiple bit-rates. Unlike previous approaches yang2023lossy ; li2024towards ; careil2024towards that start from Gaussian noise, we introduce the concept of pseudo diffusion timestep, bridging the quantization process and the forward diffusion. This idea is inspired by a classical signal processing result that quantization noise, especially at higher bit-rates, can be approximated as additive Gaussian noise bennett1948spectra . Formally, we posit that the quantized latents $\tilde{\mathbf{z}}$ can be decomposed as $\tilde{\mathbf{z}}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}$, where $\boldsymbol{\epsilon}$ is approximately Gaussian noise and statistically independent of $\mathbf{z}_{0}$, $\bar{\alpha}_{t}$ is the diffusion noise schedule, and $t$ is the pseudo diffusion timestep that depends on the bit-rate. Under the assumption that $\boldsymbol{\epsilon}$ is orthogonal to $\mathbf{z}_{0}$ in expectation, the expected cosine similarity between $\tilde{\mathbf{z}}$ and $\mathbf{z}_{0}$ satisfies $\mathbb{E}[\operatorname{sim(\tilde{\mathbf{z}},\mathbf{z}_{0})}]\approx\sqrt{\bar{\alpha}_{t}}$ (Please refer to Appendix for a full derivation). To make this relationship more consistent, we adopt a representation alignment training between $\tilde{\mathbf{z}}$ and $\mathbf{z}_{0}$. We empirically observe that this training paradigm enables their cosine similarity to converge to a fixed value for each bit-rate. Thus, by matching measured cosine similarities to the theoretical form, each bit-rate $r$ can be directly mapped to a pseudo diffusion timestep $t$. Our experiments provide strong support for this modeling, as the measured distribution of $\boldsymbol{\epsilon}$ closely follows the assumed Gaussian distribution.
@@
Figure 2: Left: Our method bridges quantization and diffusion by mapping bit-rates to pseudo diffusion timesteps. Right: BD-rates measured by DISTS ding2020image , with circle radius indicating multiply-accumulate operations (MACs). OSCAR achieves a favorable quality-efficiency tradeoff.
@@ 2.1 Neural Image Compression
Deep learning has enabled neural image compression methods to outperform traditional codecs such as JPEG wallace1991jpeg , BPG bpg , and VVC bross2021overview . Pioneering work balle2016end introduced end-to-end autoencoders for learned compression, followed by efforts balle2018variational ; minnen2020channel ; cheng2020learned ; he2022elic to jointly optimize rate and distortion through improved latent representations. Yet, distortion-based losses (e.g., MSE) often lead to distribution shifts between reconstructed and natural images blau2019rethinking , motivating research agustsson2019generative ; he2022po ; agustsson2023multi to enhance perceptual quality. HiFiC mentzer2020high explores the integration of GANs goodfellow2020generative with compression. MS-ILLM muckley2023improving improves reconstruction quality through improved discriminator design. Other methods leverage textual semantics lee2024neural , causal modeling han2024causal , or decoding-time guidance agustsson2023multi to further boost perceptual fidelity. However, the aforementioned methods typically require training separate models for each bit-rate, which results in substantial overhead.
To address the inefficiency of training separate models for each bit-rate, recent research has explored adaptive compression strategies that enable flexible bit-rate control. In particular, progressive compression  toderici2017full ; lu2021progressive ; zhang2023exploring has attracted significant interest for its ability to produce scalable bitstreams. DPICT lee2022dpict encodes latent features into trit-plane bitstreams, allowing fine-grained scalability. CTC jeon2023context extends this idea by introducing context-based trit-plane coding. Meanwhile, variable-rate approaches  chen2020variable ; wang2023evc have demonstrated effectiveness by dynamically adjusting scalar parameters or employing conditional convolutions to accommodate diverse quality levels. Despite their flexibility, these methods tend to suffer from notable quality degradation at very low bit-rates.
@@ 2.2 Diffusion Models
Diffusion models are powerful generative frameworks that transform random noise into coherent data through a multi-step denoising process ho2020denoising . Denoising Diffusion Implicit Models (DDIM) song2020denoising accelerate this process without compromising quality, influencing many subsequent advances. Latent Diffusion Models (LDMs) further improve efficiency by operating in a compressed latent space. Stable Diffusion (SD) rombach2022high advances the LDM architecture into a scalable text-to-image system, enabling flexible multimodal conditioning while maintaining strong image fidelity. In our work, we build on SD to benefit from its strong image generation capabilities.
Beyond image synthesis, diffusion models have also been explored in image compression  xu2024idempotence ; ohayon2025compressed . Early work ho2020denoising explored diffusion-based compression through reverse channel coding, with an emphasis on rate-distortion trade-offs. Follow-up studies Theis2022b extended this line by applying the technique to lossy transmission of Gaussian samples. Recent approaches further leverage large-scale pretrained diffusion models to enhance performance vonderfecht2025lossy . CDC yang2023lossy replaces the VAE decoder with a conditional diffusion model, reconstructing images via reverse diffusion conditioned on a learned content latent. PerCo careil2024towards adopts a VQ-VAE-style codebook for discrete representation and incorporates global textual descriptions to guide the iterative decoding process. DiffEIC li2024towards uses a lightweight control module to guide a frozen diffusion model with compressed content. While these models achieve strong results, their multi-step denoising process remains computationally expensive. Improving sampling efficiency is therefore essential for practical diffusion-based image compression.
@@ 3.1 Preliminaries: One-Step Latent Diffusion Model
Latent Diffusion Models (LDMs) rombach2022high operate in a lower-dimensional latent space rather than pixel space, enabling more efficient training and generation. The forward diffusion process perturbs a clean latent variable $\mathbf{z}$ by injecting Gaussian noise. At timestep $t$, the corrupted latent is given by:
@@
where the noise schedule is encoded as $\bar{\alpha}_{t}=\prod_{i=1}^{t}(1-\beta_{i})$, and $\beta_{i}$ is the variance. The reverse process aims to progressively recover the original latent through a learned denoising network ho2020denoising ; song2020denoising , typically implemented as a U-Net ronneberger2015u structure in LDMs. Standard diffusion models perform multi-step denoising via a Markov chain, where each reverse step is modeled as:
@@ 3.3 Stage 1: Learning Quantization Hyper-Encoders
Figure 4: Distribution and $Q$-plot of residual noise $\boldsymbol{\epsilon}$ in Eq. (11) before and after representation alignment. Top: red line is the standard Gaussian distribution; blue histogram is the measured distribution. Bottom: red line is the theoretical quantile line; blue dots are the measured values.
[…]
@@
where N is the number of vectors in $\mathbf{z}_{0}$ and $r$ is the desired compression bit-rate. As in VQ-VAE oord2017neural , since vector-quantization (VQ) operation is non-differentiable, the VQ loss is defined as:
@@
where $\operatorname{sg}$ is the stop-gradient operation, and $\mathbf{z}_{q}$ is the mapping of $\mathbf{z}_{0}$ to its nearest codebook entry. Following the method in esser2021taming , we adopt an exponential moving average (EMA) update for the codebook, which enhances training stability and eliminates the need for the first term in Eq. (11). Thus, the overall representation alignment objective is:
@@
We argue that representation alignment is critical for modeling compressed latents as intermediate diffusion states (We will elaborate on this in Section 3.4). To support this, we compare the measured additive noise before and after training. As shown in Fig. 4, prior to training, the residual noise exhibits no clear structure and resembles a uniform distribution within the range of 2.5 to 2.5. The corresponding quantile-quantile plots (QQ-plots) wilk1968probability deviate substantially from the theoretical Gaussian line. This indicates that the latents prior to alignment contain only unstructured noise. In contrast, after training, the residual noise closely follows a Gaussian distribution, and the QQ-plots align well with the theoretical line, indicating that the residuals become statistically well-behaved.
@@ 3.4 Stage 2: Unified Training for One-Step Reconstruction
Specifically, given the predefined bit-rate set $\mathcal{R}$, we fine-tune all the corresponding hyper-encoders and apply LoRA hu2021lora to update the diffusion model. The overall loss combines a representation alignment loss $\mathcal{L}_{\text{repa}}^{r}$, a perceptual loss $\mathcal{L}_{\text{per}}$, and an adversarial loss $\mathcal{L}_{\mathcal{G}}$:
@@
The alignment loss $\mathcal{L}_{\text{repa}}$ preserves structural consistency between the quantized and original latents. The perceptual loss $\mathcal{L}_{\text{per}}$ includes MSE and DISTS ding2020image terms. Following the practice of das2023edge , we additionally use the Sobel operator to extract the edge of the image and measure the corresponding edge-aware DISTS loss:
[…]
where $\hat{\mathbf{I}}=\mathcal{G}_{\phi,\theta}(\mathbf{z}_{0})$, $\mathcal{S}(\cdot)$ is the Sobel operator. Finally, we introduce a discriminator $\mathcal{D}_{\psi}$ and adopt a GAN loss agustsson2019generative in latent space to encourage realism of the reconstructed image: